# Variational Autoencoders for Efficient Simulation-Based Inference

### v1

@@ 3.2 Unconditional Prior VAE (UP-VAE)
A simpler alternative optimization scheme is also proposed, that employs an unconditional prior, in contrast to the earlier approach where the prior was conditioned on the observed data $\mathbf{y}$. Conditioning the prior on $\mathbf{y}$ is beneficial for amortized inference, but adds additional complexity to the model as the prior now depends on the data, potentially increasing the computational burden and complicating inference. Moreover, a data-dependent prior can become overly tailored to the training data, reducing the model’s ability to generalize to new observations. By utilizing an unconditional prior, we simplify the model architecture and promote better generalization by preventing overfitting to training data. To address this, we use Bayes’ theorem to reformulate the conditional distribution $p(\mathbf{z}\mid\boldsymbol{\theta},\mathbf{y})$ as:
$p(\mathbf{z}\mid\boldsymbol{\theta},\mathbf{y})=\frac{p(\boldsymbol{\theta},\mathbf{y}\mid\mathbf{z})p(\mathbf{z})}{p(\boldsymbol{\theta},\mathbf{y})}.$ (7)
Substituting (7) into the KL divergence formulation in (2), and denoting $D_{KL}(q_{\psi}(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta})\|p(\mathbf{z}\mid\mathbf{y},\boldsymbol{\theta}))$ as $\mathcal{D}$,
$\mathcal{D}$ $=\mathbb{E}_{\mathbf{z}\sim q_{\psi}}\bigg{[}\log q_{\psi}(\mathbf{z}\mid\boldsymbol{\theta},\mathbf{y})-\log p(\mathbf{z})\bigg{]}$ (8)
$\quad-\mathbb{E}_{\mathbf{z}\sim q_{\psi}}\bigg{[}\log p(\boldsymbol{\theta},\mathbf{y}\mid\mathbf{z})-\log p(\boldsymbol{\theta},\mathbf{y})\bigg{]}.$
Applying the chain rule to the joint probability $p(\boldsymbol{\theta},\mathbf{y})=p(\mathbf{y})\,p(\boldsymbol{\theta}\mid\mathbf{y})$ and recognizing that $\log p(\mathbf{y})$ is a constant with respect to $\boldsymbol{\theta}$, along with the non-negativity property of the KL divergence, we obtain:
$\log p(\boldsymbol{\theta}\mid\mathbf{y})$ $\geq\mathbb{E}_{\mathbf{z}\sim q_{\psi}}\bigg{[}\log p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{z})+\log p(\mathbf{y}\mid\mathbf{z})\bigg{]}$ (9)
$\quad-D_{KL}\bigg{(}q_{\psi}(\mathbf{z}\mid\boldsymbol{\theta},\mathbf{y})\,\|\,p(\mathbf{z})\bigg{)}.$
@@
$\mathcal{L}(\boldsymbol{\theta},\mathbf{y};\psi)=$ $\,D_{KL}\big{(}q_{\psi}(\mathbf{z}\mid\boldsymbol{\theta},\mathbf{y})\,\|\,p(\mathbf{z})\big{)}$ (10)
$-\mathbb{E}_{\mathbf{z}\sim q_{\psi}}\big{[}\log p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{z})\big{]}$
$-\mathbb{E}_{\mathbf{z}\sim q_{\psi}}\big{[}\log p(\mathbf{y}\mid\mathbf{z})\big{]}.$
This formulation employs two decoder networks: one for reconstructing the parameters $\boldsymbol{\theta}$, and another for reconstructing the observed data $\mathbf{y}$. The data decoder introduces an auxiliary loss, which complements the primary loss from the parameter decoder. This auxiliary objective helps the model capture latent space features that are informative for both $\boldsymbol{\theta}$ and $\mathbf{y}$ (Sheth and Kahou, 2024). However, for amortized inference, only the parameter decoder is needed, as it directly generates the parameter estimates conditioned on the observed data. Similar to CP-VAE, we approximate the prior $p(\mathbf{z})$, the encoder $q_{\psi}(\mathbf{z}\mid\boldsymbol{\theta},\mathbf{y})$, the parameter decoder $p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{z})$, and the data decoder $p(\mathbf{y}\mid\mathbf{z})$ using Gaussian distributions, simplifying the computation of intractable terms. The architecture of the model is depicted in Figure 2. The derivation of the loss function and the model’s distributional assumptions are provided in Section 2 of the Appendix.
Amortized inference using UP-VAE proceeds by first sampling $N$ latent variables $\{\mathbf{z}^{(i)}\}_{i=1}^{N}$ from the unconditional prior $p(\mathbf{z})$. These latent variables, along with the observed data $\mathbf{y}_{0}$, are then input into the trained theta decoder $p(\boldsymbol{\theta}\mid\mathbf{y}_{0},\mathbf{z}^{(i)})$, which outputs the parameters of the theta distribution from which $\boldsymbol{\theta}^{(i)}$ are sampled. Aggregating these samples empirically approximates $p(\boldsymbol{\theta}\mid\mathbf{y}_{0})$ in an efficient manner.
@@ 4 Experiments
[…]
4.1 The Two Moons Model
Figure 3: Two moons: true v/s estimated posterior (sim. budget: 10,000). (a) CP-VAE and (b) UP-VAE
The Two Moons model is a popular benchmark commonly used to evaluate the ability of SBI algorithms to recover multimodal and unconventional posterior distributions (Greenberg et al., 2019; Lueckmann et al., 2021). The challenge herein is to recover the bimodal nature of the posterior, and to capture its distinctive crescent shape. The parameter vector $\boldsymbol{\theta}\sim\mathcal{U}(-1,1)^{2}$ governs the model, and the observed data $\mathbf{x}$ is generated as:
$\mathbf{x}|\boldsymbol{\theta}=\begin{pmatrix}r\cos(a)+0.25\\ r\sin(a)\end{pmatrix}+\frac{1}{\sqrt{2}}\begin{pmatrix}-|\theta_{1}+\theta_{2}|\\ -(\theta_{1}-\theta_{2})\end{pmatrix},$ (11)
for radius $r\sim\mathcal{N}(0.1,0.01^{2})$ and angle $a\sim\mathcal{U}\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$.
Figure 3 shows the recovered posterior using (a) CP-VAE and (b) UP-VAE model formulations. Both models generate samples that align closely with the true posterior, with UP-VAE exhibiting slightly better coverage of the crescent shape. Certain discrepancies, such as outliers, may be attributed to the variational lower bound in the loss function. Tables 1 and 2 present a comparative analysis using the C2ST accuracy and MMD metrics respectively. JANA and APT perform well, accurately characterizing the bimodal posterior. CP-VAE and UP-VAE closely follow, maintaining accuracy with low variance. Notably, both CP-VAE and UP-VAE are training efficient (Table 3), while demonstrating effectiveness in handling complex, unconventional posterior distributions.
Figure 4: SIR model: true v/s estimated posterior distribution. (a) CP-VAE and (b) UP-VAE.
4.2 The Bernoulli Generalized Linear Model
The Bernoulli Generalized Linear Model (GLM) simulates neuronal spiking behavior in response to a time-varying stimulus, driven by a set of covariates (Gonzalves et al., 2020). With a discrete time domain of $T=100$ bins, the model produces observed data in the form $\mathbf{y}=[y_{i}]_{i=1}^{10}$. Each component $y_{i}$ aggregates spiking activity across time intervals, capturing both overall spiking frequency and spatial patterns of neuronal activity. The model is parameterized by a vector $\boldsymbol{\theta}\in\mathbb{R}^{10}$, consisting of a scalar bias term $\beta$ and an auxiliary vector $\mathbf{f}\in\mathbb{R}^{9}$:
$\boldsymbol{\theta}=\begin{bmatrix}\beta\\ \mathbf{f}\end{bmatrix},\quad\beta\sim\mathcal{N}(0,2),\quad\mathbf{f}\sim\mathcal{N}\left(\mathbf{0},\left(\mathbf{F}^{\top}\mathbf{F}\right)^{-1}\right),$ (12)
where the matrix $\mathbf{F}$ enforces smoothness by penalizing second-order differences in $\mathbf{f}$, defined as:
$F_{ij}=\begin{cases}1,&j=i-2,\\ -2,&j=i-1,\\ 1+\sqrt{\frac{i-1}{9}},&j=i,\\ 0,&\text{otherwise.}\end{cases},\quad 1\leq i\leq 9,$ (13)
For each time bin $i$, the binary spiking variable $z_{i}$ is sampled from a Bernoulli distribution, where the spiking probability is modeled using a logistic function applied to the linear combination of covariates $\mathbf{v}_{i}$ and the parameter vector, expressed as:
$z_{i}\sim\text{Bernoulli}\left(\sigma(\mathbf{v}_{i}^{\top}\mathbf{f}+\beta)\right),\quad\sigma(t)=\frac{1}{1+e^{-t}}.$ (14)
Here, $\mathbf{v}_{i}\in\mathbb{R}^{9}$ represents white noise inputs (covariates) drawn from $\mathcal{N}(0,\mathbf{I})$. The design matrix $\mathbf{V}=\left[\mathbf{v}_{i}\right]_{i=1}^{T}$ aggregates these inputs over time bins $T$. The model computes summary statistics $\mathbf{y}\in\mathbb{R}^{10}$, where the total number of spikes is represented by $y_{1}=\sum_{i=1}^{T}z_{i}$, and the remaining components, $y_{2},\dots,y_{10}$, are spike-triggered averages of the covariates, normalized by $y_{1}$, defined as $y_{2:10}=\frac{1}{y_{1}}\mathbf{V}\mathbf{z}$. The summary statistics provide a compact representation of the simulated spiking activity, capturing the overall frequency and the influence of covariates on the spiking process.
Figure 5 presents the posterior distributions of $\theta_{1}$ through $\theta_{10}$ for CP-VAE and UP-VAE. In both cases, the posterior distributions align well with the true posteriors, though CP-VAE exhibits a better fit as compared to UP-VAE. CP-VAE achieves the lowest C2ST accuracy (Table 1) and MMD values (Table 2), indicating a closer match to the true posterior. UP-VAE follows closely behind on the C2ST metric, while NPE is the second-best method on the MMD metric. Both CP-VAE and UP-VAE are substantially faster to train (Table 3) showcasing efficiency and scalability for this test problem involving 10 parameters.
4.3 The SIR Model
The Susceptible-Infected-Recovered (SIR) model is a widely used mathematical framework in epidemiology for modeling the spread of infectious diseases within a population. This model tracks the evolution of the number of individuals in three distinct compartments: susceptible ($S$), infected ($I$), and recovered ($R$). The dynamics of disease spread are governed by a non-linear dynamical system characterized by two key parameters: the contact rate ($\beta$) and the recovery rate ($\gamma$). The model is described by the following set of ordinary differential equations:
$\frac{d}{dt}\begin{pmatrix}S\\ I\\ R\end{pmatrix}=\begin{pmatrix}-\beta S\\ \beta S-\gamma I\\ \gamma I\end{pmatrix}\frac{I}{N},$ (15)
where $N$ is the total population size. The parameters are sampled as: $\beta\sim\text{LogNormal}(\log(0.4),0.5)$ and $\gamma\sim\text{LogNormal}(\log(1/8),0.2)$.
Figure 4 presents the posterior distributions estimated using CP-VAE and UP-VAE, respectively. Both models effectively capture complex, data-dependent relationships, resulting in accurate posterior approximations. Both CP-VAE and UP-VAE performed well on the C2ST (Table 1) and MMD metrics (Table 2), while being highly efficient (training times in Table 3).
Figure 5: Bernoulli GLM: true v/s estimated posterior distribution. (a) CP-VAE and (b) UP-VAE.
@@ 4.3 The SIR Model
Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.7770 (± 0.0630) 0.6634 (± 0.0503) 0.7467 (± 0.1036) 0.9999 (± 0.0001) 0.9999 (± 0.0000) 0.9997 (± 0.0005) 0.9990 (± 0.0015) 0.9912 (± 0.0161) 0.9838 (± 0.0157) NPE 0.5749 (± 0.0331) 0.5931 (± 0.0329) 0.5411 (± 0.0195) 0.9121 (± 0.0338) 0.8988 (± 0.0306) 0.8898 (± 0.0565) 0.8378 (± 0.0396) 0.9122 (± 0.0207) 0.9405 (± 0.0125) JANA 0.5206 (± 0.0053) 0.5151 (± 0.0055) 0.5123 (± 0.0032) 0.9678 (± 0.0035) 0.9692 (± 0.0026) 0.9589 (± 0.0034) 0.8697 (± 0.0149) 0.9175 (± 0.0059) 0.9343 (± 0.0055) APT 0.5189 (± 0.0080) 0.5202 (± 0.0114) 0.5147 (± 0.0125) 0.8404 (± 0.0107) 0.8328 (± 0.0156) 0.8353 (± 0.0164) 0.9433 (± 0.0076) 0.9434 (± 0.0061) 0.9457 (± 0.0068) CP-VAE 0.6443 (± 0.0288) 0.6072 (± 0.0108) 0.6346 (± 0.0303) 0.6955 (± 0.0239) 0.6697 (± 0.0361) 0.6427 (± 0.0177) 0.8524 (± 0.0719) 0.8976 (± 0.0121) 0.8857 (± 0.0330) UP-VAE 0.6510 (± 0.0310) 0.6565 (± 0.0364) 0.6212 (± 0.0252) 0.8270 (± 0.0249) 0.8311 (± 0.0216) 0.7819 (± 0.0692) 0.8390 (± 0.0378) 0.8731 (± 0.0482) 0.8893 (± 0.0317)

### v2

@@ 2.2 Unconditional Prior VAE (UP-VAE)
A simpler alternative optimization scheme is also proposed, that employs an unconditional prior, in contrast to the earlier approach where the prior was conditioned on the observed data $\mathbf{y}$. Conditioning the prior on $\mathbf{y}$ is beneficial for amortized inference, but adds additional complexity to the model as the prior now depends on the data, potentially increasing the computational burden and complicating inference. Moreover, a data-dependent prior can become overly tailored to the training data, reducing the model’s ability to generalize to new observations. By utilizing an unconditional prior, we simplify the model architecture and promote better generalization by preventing overfitting to training data. To address this, we use Bayes’ theorem to reformulate the conditional distribution $p(\mathbf{z}\mid\bm{\theta},\mathbf{y})$ as:
$p(\mathbf{z}\mid\bm{\theta},\mathbf{y})=\frac{p(\bm{\theta},\mathbf{y}\mid\mathbf{z})p(\mathbf{z})}{p(\bm{\theta},\mathbf{y})}.$ (8)
Substituting (8) into the KL divergence formulation in (2), and denoting $D_{KL}(q_{\psi}(\mathbf{z}\mid\mathbf{y},\bm{\theta})\|p(\mathbf{z}\mid\mathbf{y},\bm{\theta}))$ as $\mathcal{D}$,
$\mathcal{D}$ $=\mathbb{E}_{\mathbf{z}\sim q_{\psi}}\bigg[\log q_{\psi}(\mathbf{z}\mid\bm{\theta},\mathbf{y})-\log p(\mathbf{z})\bigg]-\mathbb{E}_{\mathbf{z}\sim q_{\psi}}\bigg[\log p(\bm{\theta},\mathbf{y}\mid\mathbf{z})-\log p(\bm{\theta},\mathbf{y})\bigg].$ (9)
Applying the chain rule to the joint probability $p(\bm{\theta},\mathbf{y})=p(\mathbf{y})\,p(\bm{\theta}\mid\mathbf{y})$ and recognizing that $\log p(\mathbf{y})$ is a constant with respect to $\bm{\theta}$, along with the non-negativity property of the KL divergence, we obtain:
$\log p(\bm{\theta}\mid\mathbf{y})$ $\geq\mathbb{E}_{\mathbf{z}\sim q_{\psi}}\bigg[\log p(\bm{\theta}\mid\mathbf{y},\mathbf{z})+\log p(\mathbf{y}\mid\mathbf{z})\bigg]-D_{KL}\bigg(q_{\psi}(\mathbf{z}\mid\bm{\theta},\mathbf{y})\,\|\,p(\mathbf{z})\bigg).$ (10)
@@
$\mathcal{L}(\bm{\theta},\mathbf{y};\psi)=$ $\,D_{KL}\big(q_{\psi}(\mathbf{z}\mid\bm{\theta},\mathbf{y})\,\|\,p(\mathbf{z})\big)-\mathbb{E}_{\mathbf{z}\sim q_{\psi}}\big[\log p(\bm{\theta}\mid\mathbf{y},\mathbf{z})\big]-\mathbb{E}_{\mathbf{z}\sim q_{\psi}}\big[\log p(\mathbf{y}\mid\mathbf{z})\big].$ (11)
This formulation employs two decoder networks: one for reconstructing the parameters $\bm{\theta}$ and another for reconstructing the observed data $\mathbf{y}$. The data decoder introduces an auxiliary loss, complementing the primary loss from the parameter decoder. This auxiliary objective ensures the latent space captures features informative for both $\bm{\theta}$ and $\mathbf{y}$ [Sheth and Kahou, 2024]. For amortized inference, only the parameter decoder is required, as it directly generates parameter estimates conditioned on $\mathbf{y}$. Similar to the previous approach, we approximate the prior $p(\mathbf{z})\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, the encoder $q_{\psi}(\mathbf{z}\mid\bm{\theta},\mathbf{y})\sim\mathcal{N}(\bm{\mu}_{q},\bm{\Sigma}_{q})$, the parameter decoder $p(\bm{\theta}\mid\mathbf{y},\mathbf{z})\sim\mathcal{N}(\bm{\mu}_{\theta},\bm{\Sigma}_{\theta})$, and the data decoder $p(\mathbf{y}\mid\mathbf{z})\sim\mathcal{N}(\bm{\mu}_{\mathbf{y}},\bm{\Sigma}_{\mathbf{y}})$ using Gaussian distributions with diagonal covariance matrices, simplifying intractable computations. The model architecture is illustrated in Figure 2. The final loss is:
$\mathcal{L}_{\text{UP-VAE}}$ $=\frac{1}{2}\left(-\log|\bm{\Sigma}_{q}|-k_{\mathbf{z}}+\mathrm{tr}(\bm{\Sigma}_{q})+\bm{\mu}_{q}^{\top}\bm{\mu}_{q}\right)+\sum_{i=1}^{k_{\theta}}\log\sigma_{\theta_{i}}+\frac{1}{2N}\sum_{j=1}^{N}\sum_{i=1}^{k_{\theta}}\frac{(\theta_{i,j}-\mu_{\theta_{i}})^{2}}{\sigma_{\theta_{i}}^{2}}$ (12)
$\quad+\sum_{i=1}^{k_{\mathbf{y}}}\log\sigma_{\mathbf{y}_{i}}+\frac{1}{2N}\sum_{j=1}^{N}\sum_{i=1}^{k_{\mathbf{y}}}\frac{(y_{i,j}-\mu_{\mathbf{y}_{i}})^{2}}{\sigma_{\mathbf{y}_{i}}^{2}},$
where $k_{\mathbf{z}}$, $k_{\theta}$, and $k_{\mathbf{y}}$ denote the dimensionalities of $\mathbf{z}$, $\bm{\theta}$, and $\mathbf{y}$, respectively.
Amortized inference using UP-VAE proceeds by first sampling $N$ latent variables $\{\mathbf{z}^{(i)}\}_{i=1}^{N}$ from the unconditional prior $p(\mathbf{z})$. These latent variables, along with the observed data $\mathbf{y}_{0}$, are then input into the trained theta decoder $p(\bm{\theta}\mid\mathbf{y}_{0},\mathbf{z}^{(i)})$, which outputs the parameters of the theta distribution from which $\bm{\theta}^{(i)}$ are sampled. Aggregating these samples empirically approximates $p(\bm{\theta}\mid\mathbf{y}_{0})$ in an efficient manner.
@@ 3 Experiments
We evaluate the proposed approach on ten benchmark problems from the sbibm suite [Lueckmann et al., 2021], each presenting unique challenges for simulation-based inference. For each test problem, the simulation budgets is varied in {10,000, 20,000, 30,000}. We then estimate the posterior distributions by generating 10,000 samples from fixed observed data across five independent runs. To evaluate the accuracy of the posterior approximations, we compare them against reference posteriors obtained via MCMC sampling, using the Maximum Mean Discrepancy (MMD) and the Classifier Two-Sample Test (C2ST) [Friedman, 2003] evaluation metrics. For consistency and reproducibility, we utilized pre-generated observed data and reference posteriors from the sbibm library [Lueckmann et al., 2021]. The C2ST metric computations follow [Lueckmann et al., 2021] – a C2ST score close to 0.5 indicates that the two distributions are indistinguishable, whereas a score significantly higher than 0.5 suggests notable differences. The MMD calculation were performed using the BayesFlow framework [Radev et al., 2023b]. All experiments were performed on a machine with 2 $\times$ 16 core Intel(R) Xeon(R) Gold 6226R CPU, 576GB RAM and a NVIDIA Tesla T4 GPU (16 GB RAM).
3.1 SBI Benchmark Problems
This section outlines distinctive challenges posed by each SBI benchmarks. The Two Moons problem tests methods ability to capture bimodal crescent-shaped posteriors. In the Gaussian Mixture Model, the main challenge is identifying the shared mean between two Gaussian distributions with different variances. SLCP introduces complex parameter dependencies (uniform, Gaussian, multimodal, and skewed distributions), while its distractors variant adds uninformative dimensions that obscure inference. Gaussian Linear examines posterior estimation under conjugate priors, whereas Gaussian Linear Uniform extends this challenge to non-conjugate uniform priors. In Bernoulli GLM , the challenge involves estimating posteriors from Bernoulli observations with Gaussian priors with the Raw variant omitting sufficient statistics to increase complexity. The SIR epidemic model aims to infer contact and recovery rates from noisy infection data over time. Finally, the Lotka-Volterra model is challenging due to its nonlinear dynamics and potential chaotic behavior.
(a) CP-VAE
(b) UP-VAE
Figure 3: Two moons: true v/s estimated posterior (sim. budget: 30,000). (a) CP-VAE and (b) UP-VAE.
(a) CP-VAE
(b) UP-VAE
Figure 4: SLCP : true v/s estimated posterior distribution (sim. budget: 30,000). (a) CP-VAE and (b) UP-VAE.
3.2 Posterior Visualization on Multimodal Benchmarks
In this section, we present posterior plots for two challenging benchmark problems with multimodal and distinctive shapes. Figure 3 shows that both CP-VAE and UP-VAE successfully capture the complex shape and multimodal distribution of the data of the two moons model. Additionally, Figure 4 demonstrates their ability to model the true posterior’s shape across dimensions, despite varying distributions. These results highlight the models’ robustness in handling complex posterior structures.
@@ 3.2 Posterior Visualization on Multimodal Benchmarks
(a) Two Moons, Bernoulli GLM, and SIR
Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.7770 $\pm$0.0630 0.6634 $\pm$0.0503 0.7467 $\pm$0.1036 0.9999 $\pm$0.0001 0.9999 $\pm$0.0000 0.9997 $\pm$0.0005 0.9990 $\pm$0.0015 0.9912 $\pm$0.0161 0.9838 $\pm$0.0157 NPE 0.5749 $\pm$0.0331 0.5931 $\pm$0.0329 0.5411 $\pm$0.0195 0.9121 $\pm$0.0338 0.8988 $\pm$0.0306 0.8898 $\pm$0.0565 0.8378 $\pm$0.0396 0.9122 $\pm$0.0207 0.9405 $\pm$0.0125 APT 0.5189 $\pm$0.0080 0.5202 $\pm$0.0114 0.5147 $\pm$0.0125 0.8404 $\pm$0.0107 0.8328 $\pm$0.0156 0.8353 $\pm$0.0164 0.9433 $\pm$0.0076 0.9434 $\pm$0.0061 0.9457 $\pm$0.0068 JANA 0.5206 $\pm$0.0053 0.5151 $\pm$0.0055 0.5123 $\pm$0.0032 0.9678 $\pm$0.0035 0.9692 $\pm$0.0026 0.9589 $\pm$0.0034 0.8697 $\pm$0.0149 0.9175 $\pm$0.0059 0.9343 $\pm$0.0055 Simformer 0.5273 $\pm$0.0229 0.5190 $\pm$0.0174 0.5098 $\pm$0.0084 0.6584 $\pm$0.0381 0.6131 $\pm$0.0225 0.6154 $\pm$0.0265 0.9378 $\pm$0.0053 0.9520 $\pm$0.0064 0.9478 $\pm$0.0049 CP-VAE 0.6443 $\pm$0.0288 0.6072 $\pm$0.0108 0.6346 $\pm$0.0303 0.6955 $\pm$0.0239 0.6697 $\pm$0.0361 0.6427 $\pm$0.0177 0.8524 $\pm$0.0719 0.8976 $\pm$0.0121 0.8857 $\pm$0.0330 UP-VAE 0.6510 $\pm$0.0310 0.6565 $\pm$0.0364 0.6212 $\pm$0.0252 0.8270 $\pm$0.0249 0.8311 $\pm$0.0216 0.7819 $\pm$0.0692 0.8390 $\pm$0.0378 0.8731 $\pm$0.0482 0.8893 $\pm$0.0317
(b) SLCP Distractors, Bernoulli GLM Raw, and Lotka Volterra
Method SLCP Distractors Bernoulli GLM Raw Lotka Volterra Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.9995 $\pm$0.0007 0.9978 $\pm$0.0016 0.9994 $\pm$0.0004 0.9998 $\pm$0.0001 0.9999 $\pm$0.0001 0.9999 $\pm$0.0000 0.9999 $\pm$0.0001 0.9999 $\pm$0.0000 1.0000 $\pm$0.0000 NPE 0.9714 $\pm$0.0144 0.9468 $\pm$0.0157 0.9359 $\pm$0.0127 0.9354 $\pm$0.0093 0.9114 $\pm$0.0082 0.9065 $\pm$0.0105 0.9999 $\pm$0.0001 0.9999 $\pm$0.0000 0.9999 $\pm$0.0000 JANA 0.9987 $\pm$0.0010 0.9945 $\pm$0.0138 0.9933 $\pm$0.0654 0.9997 $\pm$0.0007 1.0000 $\pm$0.0000 1.0000 $\pm$0.0000 0.9999 $\pm$0.0000 0.9998 $\pm$0.0001 0.9998 $\pm$0.0001 APT 0.9823 $\pm$0.0055 0.9451 $\pm$0.0158 0.9130 $\pm$0.0301 0.8455 $\pm$0.0116 0.8596 $\pm$0.0152 0.8461 $\pm$0.0135 0.9999 $\pm$0.0001 0.9999 $\pm$0.0000 0.9999 $\pm$0.0001 Simformer 0.9198 $\pm$0.0451 0.8814 $\pm$0.0433 0.8717 $\pm$0.0453 0.9814 $\pm$0.0106 0.8600 $\pm$0.0452 0.7750 $\pm$0.0473 0.9999 $\pm$0.0003 0.9999 $\pm$0.0012 1.000 $\pm$0.0002 CP-VAE 0.9956 $\pm$0.0037 0.9880 $\pm$0.0057 0.9746 $\pm$0.0075 0.9491 $\pm$0.0078 0.9171 $\pm$0.0258 0.9307 $\pm$0.0000 0.9999 $\pm$0.0001 0.9999 $\pm$0.0001 0.9999 $\pm$0.0001 UP-VAE 0.9838 $\pm$0.0036 0.9798 $\pm$0.0023 0.9840 $\pm$0.0053 0.8674 $\pm$0.0493 0.8790 $\pm$0.0363 0.8536 $\pm$0.0460 0.9999 $\pm$0.0001 0.9999 $\pm$0.0001 0.9999 $\pm$0.0001
(c) Gaussian Mixture, Gaussian Linear, Gaussian Linear Uniform, and SLCP
Method Gaussian Mixture Gaussian Linear Gaussian Linear Uniform SLCP Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.7474 $\pm$0.0364 0.6977 $\pm$0.0260 0.7234 $\pm$0.0599 0.9930 $\pm$0.0038 0.9916 $\pm$0.0046 0.9929 $\pm$0.0059 0.9989 $\pm$0.0004 0.9976 $\pm$0.0016 0.9988 $\pm$0.0010 0.9687 $\pm$0.0286 0.9436 $\pm$0.0203 0.9723 $\pm$0.0140 NPE 0.6118 $\pm$0.0155 0.5830 $\pm$0.0240 0.5818 $\pm$0.0158 0.5472 $\pm$0.0076 0.5238 $\pm$0.0112 0.5266 $\pm$0.0037 0.5562 $\pm$0.0147 0.5292 $\pm$0.0068 0.5256 $\pm$0.0066 0.8996 $\pm$0.0206 0.8502 $\pm$0.0174 0.8546 $\pm$0.0213 JANA 0.6146 $\pm$0.0066 0.5955 $\pm$0.0026 0.5854 $\pm$0.0037 0.9905 $\pm$0.0028 0.9903 $\pm$0.0003 0.9904 $\pm$0.0009 0.9845 $\pm$0.0027 0.9854 $\pm$0.0021 0.9879 $\pm$0.0011 0.8707 $\pm$0.0345 0.8421 $\pm$0.0292 0.7791 $\pm$0.0232 APT 0.5457 $\pm$0.0058 0.5425 $\pm$0.0124 0.5467 $\pm$0.0050 0.5367 $\pm$0.0038 0.5293 $\pm$0.0052 0.5258 $\pm$0.0047 0.5401 $\pm$0.0115 0.5302 $\pm$0.0069 0.5264 $\pm$0.0075 0.8164 $\pm$0.0295 0.6882 $\pm$0.0342 0.6473 $\pm$0.0343 Simformer 0.5166 $\pm$0.0210 0.5100 $\pm$0.0090 0.5126 $\pm$0.0070 0.6130 $\pm$0.0281 0.6226 $\pm$0.0398 0.5866 $\pm$0.0430 0.6501 $\pm$0.0519 0.6207 $\pm$0.0519 0.5780 $\pm$0.0211 0.7661 $\pm$0.0685 0.6201 $\pm$0.0490 0.7192 $\pm$0.0721 CP-VAE 0.7328 $\pm$0.0347 0.7492 $\pm$0.0496 0.7482 $\pm$0.0228 0.6309 $\pm$0.0357 0.6003 $\pm$0.0165 0.5851 $\pm$0.0355 0.7854 $\pm$0.0269 0.7169 $\pm$0.0332 0.6886 $\pm$0.0637 0.9279 $\pm$0.0040 0.8987 $\pm$0.0164 0.8582 $\pm$0.0148 UP-VAE 0.7977 $\pm$0.0390 0.8134 $\pm$0.0395 0.8032 $\pm$0.0418 0.7704 $\pm$0.0461 0.6037 $\pm$0.0318 0.7530 $\pm$0.0060 0.8024 $\pm$0.0156 0.7898 $\pm$0.0137 0.7549 $\pm$0.0276 0.9578 $\pm$0.0168 0.9417 $\pm$0.0114 0.9409 $\pm$0.0162
@@
(a) Two Moons, Bernoulli GLM, and SIR
Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.1213 $\pm$0.0650 0.0582 $\pm$0.0122 0.1432 $\pm$0.1161 3.6381 $\pm$0.5428 4.8299 $\pm$0.6825 3.9965 $\pm$1.8510 2.9470 $\pm$0.4891 2.7502 $\pm$0.7268 2.1466 $\pm$0.5625 NPE 0.0254 $\pm$0.0025 0.0241 $\pm$0.0099 0.0125 $\pm$0.0045 1.0371 $\pm$0.1830 0.9379 $\pm$0.1001 0.8675 $\pm$0.1945 0.6611 $\pm$0.2236 1.0312 $\pm$0.1255 1.1903 $\pm$0.1007 JANA 0.0060 $\pm$0.0013 0.0068 $\pm$0.0017 0.0047 $\pm$0.0009 3.3232 $\pm$0.0932 3.1488 $\pm$0.0280 2.9010 $\pm$0.0878 0.8495 $\pm$0.1254 1.1077 $\pm$0.0336 1.1421 $\pm$0.0392 APT 0.0073 $\pm$0.0030 0.0069 $\pm$0.0030 0.0037 $\pm$0.0012 0.7383 $\pm$0.0853 0.4944 $\pm$0.0663 0.5808 $\pm$0.0539 1.1003 $\pm$0.0712 1.1093 $\pm$0.0311 1.1291 $\pm$0.0481 Simformer 0.0005 $\pm$0.0003 0.0001 $\pm$0.0001 0.0003 $\pm$0.0004 0.0191 $\pm$0.0098 0.0156 $\pm$0.0087 0.0116 $\pm$0.0061 0.0018 $\pm$0.0028 0.0016 $\pm$0.0042 0.0016 $\pm$0.0085 CP-VAE 0.0382 $\pm$0.0110 0.0369 $\pm$0.0088 0.0276 $\pm$0.0050 0.1966 $\pm$0.1375 0.1557 $\pm$0.0663 0.0942 $\pm$0.0547 0.8792 $\pm$0.3327 1.0158 $\pm$0.0566 0.9990 $\pm$0.1923 UP-VAE 0.0736 $\pm$0.0097 0.0795 $\pm$0.0151 0.0544 $\pm$0.0150 1.5420 $\pm$0.4930 1.4748 $\pm$0.2623 1.1160 $\pm$0.7496 0.5168 $\pm$0.1193 0.6884 $\pm$0.2044 0.6487 $\pm$0.1087
(b) SLCP Distractors, Bernoulli GLM Raw, and Lotka Volterra